A car of mass 1050 kg accelerates uniformly at 2.4 ms-2 from an initial speed of 10 ms-1. It travels a distance of 540 m before it reaches its top speed.

a. Calculate its top speed.

b. The total resistive force on the car at its top speed is 3000 N. What is the power of the engine in Watts?

c. Given that 1 horsepower = 745.7 Watts, what is the horsepower rating of the car engine (to two significant figures)?

d. What is the kinetic energy of the car at its top speed?

e. Please calculate its momentum.


1
Expert's answer
2022-01-31T07:34:54-0500

a)

"a=\frac{v_2^2-v_1^2}{2s},\implies v_2=\sqrt{v_1^2+2as}=51.9~\frac ms ,"

b)

"W=Fv=0.15~MW,"

c)

"W_h=\frac Wh=208.73~hp,"

d)

"E_k=\frac{mv_2^2}2=1.4~MJ,"

e)

"p=mv_2=54.5\cdot 10^3~kg\cdot \frac ms."